Equilibrium Constant

Q1

for both of the following equilibria, write the expression for the equilibrium constant Kc and state it’s units:

i) H2(g) + I2 (g) —-> <—— 2HI (g) (2 marks)

ii) 2SO2 (g) + O2(g) ——> <—— 2SO2 (g) (2 marks)

Q2

the reaction for the formation of hydrogen bromine does not go to completion but reaches an equilibrium, as shown:

H2(g) + Br2 —-> <—— 2HBr(g)

a mixture of 1.9 mol of H2 and 1.9 of Br was prepared and allowed to reach equilibrium in a closed vessel on 250 cm3 capacity. The resulting equilibrium mixture was found to contain 3.0 moles of HBr. Calculate the value of K. (3 marks please detailed answers with explanation🙏🏾🙏🏾🙏🏾🙏🏾)


1
Expert's answer
2022-01-18T18:11:02-0500

Q1

(i) "k_c=\frac{[HI]^2}{[H_{2}][I_{2}]}"


Units Mole L-1


(ii) "k_c = \frac{[SO_2]^2}{[SO_2]^2[O_2]}"


Units Mole-1L


Q2


"\begin{matrix}
& H_2+ &Br_2&\rightleftharpoons&2HBr \\
 I&1.9&1.9& & 0\\
C&x&x&&2x\\
E&1.9-x&1.9-x&&2x
\end{matrix}"



But "2x=3\implies\>x=1.5"

"[H_2]=1.9-1 .5"

"=0.4mol\/250cm^3"

"=1.6mol\/L"


"[Br_2]=1.6mol\/L"

"[HBr]=\frac{3×1000}{250}=12mol/L"


"K_c=\frac{[HBr]^2}{[H_2][Br_2]}=\frac{12^2}{1.6×1.6}=56.25"